Ticket Number: HEA-NOC /20020516-01 Ticket Status: OPEN
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: EPA
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: 20020516 11:28:43 UTC Problem Start:
Ticket Closed: Problem End:
EPA are upgrading their Firewall on Friday the 16th of May.
All connectivity to the EPA.
Mail will be backed up by heanet
Time to Fix:
Outage expected for all of tomorrow, starting in mid morning